OpenStudy (anonymous):

Solve ln(4x-2x^(2))-ln(x)=ln(2x-1)

OpenStudy (anonymous):

ln(a) - ln(b) = ln(a/b)

OpenStudy (anonymous):

can you possibly show me the steps of this problem

OpenStudy (anonymous):

\[\ln(4x - 2x^2) - \ln(x) = \ln(2x-1)\] \[\ln((4x-2x^2)/x) = \ln(2x-1)\] \[\ln(4-2x) = \ln(2x-1)\] \[4-2x = 2x-1\] \[4 = 4x -1\] \[ 5 = 4x\] \[5/4 = x\] checking if x=5/4 satisfies the equation \[\ln (4 - 2(5/4)) = \ln(2 (5/4) - 1)\] \[\ln(4 - (5/2)) = \ln(5/2 - 1)\] \[\ln (3/2) = \ln (3/2)\] true statement. therefore your solution is x = 5/4

OpenStudy (anonymous):

thank you so much